A rock is thrown from the top of a 36 m high cliff with an initial speed of 12 ms–1 at an angle of 52° to the horizon. How long will it take the rock to reach the bottom of the cliff?


1
Expert's answer
2022-04-13T09:20:47-0400

"t=\sqrt\frac{2s}g+\frac{v\sin\alpha}g=14~s."